# Adaptive Gradient Calibration for Single-Positive Multi-Label Learning in Remote Sensing Image Scene Classification

### v1

@@ IV-C3 AID
Table V displays the performance of all methods on the AID dataset with Random SPML noise. This dataset, benefiting from ImageNet pre-training and richer spatial information, generally shows higher absolute performance metrics across the board compared to reBEN, despite its generally higher flip rates $\beta_{c}$. Consistent with its overall very good performance, AdaGC emerges at the top of the ranking for most metrics. It achieves the best mAP score, and the second best mF1, mrecall, coverage, and rankloss scores. Notably, Dist-PU shows outstanding performance, on par with AdaGC’s mF1, and achieves the best OA, coverage, and rankloss scores. Additionally, EPR shows the best mF1 score. Nevertheless, the top performance of Dist-PU and EPR can be explained in terms of prior information available. Indeed, both methods exploit some a-priori information regarding the GT multi-labels, which is not the case for the proposed AdaGC. Dist-PU assumes that the knowledge of the prior distributions of all classes is available, and exploits this information to match it with the class priors of the predicted labels, while EPR uses the knowledge of the expected number of positive labels per image. This information is particularly effective on the AID dataset, as it shows, in general, a higher number of labels for each image, as shown in Fig. 4. This makes EPR exhibit the best mrecall among all methods, on par with GT training, which is high enough to counterbalance the worst mprecision across the board, which highlights a label overestimation tendency.
@@ IV-D Ablation Experiments
IV-D1 Early Learning Detection
(a) Random-mAP
(b) Random-mF1
(c) Dominant-mAP
(d) Dominant-mF1
Figure 5: Validation accuracies (mAP and mF1) versus training time (epoch) obtained with predefined warm-up length (10, 20, 30) and our proposed early learning detection strategy (AdaGC).
We evaluate the effectiveness of early learning detection for adaptively triggering GC by comparing AdaGC with variants that use manually predefined trigger points. Fig. 5 plots their true validation accuracies (with respect to the GT labels) over training time. As shown, the proposed adaptive detection strategy successfully activates GC at an appropriate point, resulting in the most significant performance improvements. In contrast, triggering GC too early or too late leads to suboptimal results. Notably, activating GC too late leads to even worse performance, highlighting the detrimental effect of overfitting to label noise. Furthermore, Table VI reports the final test accuracies. AdaGC achieves competitive or comparable mAP and mF1 scores compared to those obtained by using fixed warm-up lengths, indicating the robustness of the proposed early learning detection strategy.
TABLE VI: Accuracies (%) on the Test Set Obtained With Different GC Trigger Settings After 70 Epochs
@@ IV-D1 Early Learning Detection
IV-D2 Teacher and Student Models
We examine the roles of the student and teacher models in two modules, i.e., early learning detection and pseudo-label generation.
We first illustrate the smoothness of the noisy validation accuracies produced by the student and teacher models in Fig. 6. The accuracy curves from both models exhibit similar trends. Nevertheless, those from the teacher models are noticeably smoother than those from the student models. This makes teacher models more suitable for early learning detection. In contrast, using student models for detection is unstable, as the detection is more likely to fall into a local maximum, leading to suboptimal results.
@@ IV-D2 Teacher and Student Models
TABLE VII: Accuracies (%) on the Test Set Obtained With Pseudo Labels
Generated by Different Models, Where T and S Represent
Teacher and Student Models, Respectively
Random Dominant
@@
AdaGC (T+S, $\gamma=0.5$) 68.18 64.65 60.22 58.78
Only T ($\gamma=1$) 61.74 63.49 42.01 46.40
Only S ($\gamma=0$) 67.93 61.52 61.11 57.42
T+S (S w/o EMA, $\gamma=0.5$) 53.78 61.94 27.50 35.71
We then investigate the roles of the student and teacher models in pseudo-label generation. As shown in Table VII, combining predictions from both models yields the most robust performance. In the Random noise setting, relying solely on either the student or the teacher model fails to achieve optimal mAP and mF1 scores. In the Dominant noise setting, using only the student model produces more stable results than using only the teacher model. This is partly due to the teacher model’s relatively slow update rate, which lags behind the student model’s more responsive prediction updates. As a result, the teacher may fail to adapt quickly enough to changing training dynamics, making it more vulnerable to overfitting to the dominant noisy labels. In such scenarios, swift and timely pseudo-label updates are essential for maintaining performance and robustness. Nevertheless, combining predictions from both models offers a balance between stability and adaptability, ensuring robust performance across different types of label noise.
Finally, we evaluate the role of applying EMA to student predictions by replacing them with the student model’s direct online predictions. The results, reported in the last row of Table VII, show a significant performance drop in both cases, with a more sever decrease in the Dominant setting. This highlights the importance of using EMA to stabilize the student predictions and ensure higher-quality pseudo-labels in GC.
@@ IV-D3 Role of Mixup
We remove Mixup from AdaGC and report the results in Table VIII to assess its impact. As shown, the absence of Mixup leads to a consistent drop in performance across both noise settings, particularly in the Random case, where mAP and mF1 decrease by approximately 2%. These findings suggest that Mixup enhances generalization, especially under random noise, by encouraging smoother decision boundaries. Although its effect is less pronounced under Dominant noise, it still provides measurable benefits, underscoring its complementary role in improving the robustness of AdaGC.
TABLE VIII: Accuracies (%) on the Test Set Obtained With and Without Mixup

### v2

@@ IV-C3 AID-Random
Table V reports the results on the AID dataset under Random SPML noise. Owing to ImageNet pre-training and richer spatial information, this dataset yields higher absolute performance than reBEN, despite larger flip rates $\beta_{c}$. AdaGC shows the best performance across all metrics. It achieves the highest mAP and the second-best mF1, mrecall, coverage, and rankloss scores. Dist-PU also shows excellent results, matching AdaGC in mF1 while achieving the best OA, coverage, and rankloss. Meanwhile, EPR reaches the highest mF1 overall. The strong outcomes of Dist-PU and EPR are largely attributable to their use of prior information about the underlying GT label distributions, which is not available to AdaGC. Dist-PU assumes knowledge of class-prior distributions and aligns predicted label frequencies accordingly, whereas EPR relies on the expected number of positive labels per image. These priors are particularly beneficial for AID, which contains more labels per image on average (Fig. 4). As a result, EPR achieves the highest mrecall, although this comes at the cost of the lowest mprecision, reflecting its tendency to overestimate positives. Still, RCML achieves some comparable results to other considered methods, whereas the improvement brought by LCR remains limited.
IV-C4 AID-Manual
TABLE VI: Test Performance Comparison of Different Methods on the AID-Manual Dataset. The Best Average Metric Values Are Reported in Bold, Second Bests Are in Italic. The Related Standard Deviations Are Reported in Brackets
$\textbf{mAP}(\%)$ Coverage $\downarrow$ Rankloss (%) $\downarrow$ OA (%) $\uparrow$ mF1 (%) $\uparrow$ mprecision (%) $\uparrow$ mrecall (%) $\uparrow$
Ground-truth Training GT w/ BCE $89.01(0.32)$ 5.834 (0.011) 1.657 (0.061) 94.29 (0.19) 83.12 (0.66) 88.87 (0.38) 79.58 (1.28)
IUN $82.71(0.60)$ 6.638 (0.085) 3.068 (0.191) 91.47 (0.16) 74.83 (0.42) 94.18 (3.95) 65.53 (0.89)
Baseline Methods AN $67.66(0.73)$ 8.589 (0.140) 11.276 (0.283) 71.30 (0.08) 29.01 (0.59) 83.94 (1.69) 18.83 (0.56)
AN-LS $66.75(0.62)$ 11.556 (0.061) 23.300 (0.640) 71.34 (0.17) 28.41 (1.81) 84.61 (1.94) 18.32 (1.19)
WAN $70.28(0.50)$ 7.721 (0.103) 8.364 (0.276) 75.35 (0.29) 45.28 (1.77) 88.30 (4.71) 32.21 (1.68)
EPRa $71.92(1.36)$ 7.251 (0.118) 8.766 (0.358) 84.89 (0.50) 71.30 (1.06) 63.86 (1.49) 83.84 (0.56)
General Label Noise Methods ELR $76.99(0.51)$ 7.278 (0.107) 5.353 (0.232) 88.26 (0.31) 69.09 (0.63) 77.43 (0.47) 68.55 (1.33)
$L$-Cp $69.65(0.54)$ 7.854 (0.073) 8.982 (0.470) 70.87 (0.07) 23.14 (0.23) 80.08 (3.81) 14.94 (0.28)
$L$-Ct $72.46(0.11)$ 7.564 (0.104) 7.788 (0.525) 76.04 (0.12) 42.70 (0.69) 85.52 (2.13) 30.93 (0.24)
$L$-R $73.89(0.29)$ 7.743 (0.007) 6.560 (0.094) 83.95 (0.05) 63.51 (1.55) 80.53 (1.29) 54.60 (1.90)
BoostLU $71.35(1.77)$ 8.441 (0.061) 8.976 (0.247) 75.61 (0.06) 42.58 (0.40) 83.31 (6.37) 31.62 (0.56)
PUL Methods Dist-PUa $74.90(1.88)$ 6.574 (0.021) 3.652 (0.085) 90.23 (0.34) 66.77 (1.16) 80.98 (0.23) 62.70 (1.32)
SPML Methods ROLE $64.36(1.99)$ 13.253 (0.101) 17.429 (0.973) 76.41 (0.64) 49.11 (2.03) 86.08 (4.40) 36.30 (2.01)
GR $67.18(0.65)$ 11.042 (0.191) 21.481 (0.353) 71.63 (0.14) 29.88 (1.34) 86.64 (4.63) 19.93 (1.17)
LAC $65.80(0.81)$ 7.878 (0.066) 9.395 (0.259) 72.64 (0.16) 31.72 (1.58) 80.81 (4.74) 21.03 (1.04)
MIME $65.39(1.07)$ 7.240 (0.147) 6.333 (0.424) 73.10 (0.14) 32.62 (0.89) 84.73 (6.70) 21.70 (0.65)
RS-specific Methods RCML $72.87(0.57)$ 7.453 (0.019) 7.393 (0.266) 80.46 (1.44) 52.99 (3.67) 85.83 (2.34) 43.08 (5.79)
LCR $66.61(0,96)$ 8.322 (0.078) 10.669 (0.472) 71.79 (0.22) 31.83 (1.46) 85.52 (1.57) 21.03 (1.02)
Proposed AdaGC $79.07(0.91)$ 6.673 (0.103) 4.620 (0.558) 82.66 (1.01) 71.95 (0.94) 68.25 (2.35) 82.61 (1.28)
Table VI reports the results on the AID dataset using Manual SPML labels to further evaluate AdaGC’s practical effectiveness. Benefiting from human-induced bias in single-positive labels, most methods achieve higher performance compared with AID-Random. In this setting, AdaGC remains highly effective, getting the highest mAP and mF1, and ranking among the top methods across nearly all metrics. Dist-PU again achieves the best OA, coverage, and rankloss by leveraging extra prior information. RS-specific methods show similar trends as before. RCML provides moderate improvements over other SPML or label-noise baselines, while LCR contributes little due to the fixed class-correlation assumptions. Overall, the AID-Manual results confirm that higher-quality single-positive labels enhance learning, and AdaGC consistently delivers robust performance without relying on additional priors.
@@ IV-D Ablation Experiments
IV-D1 Impact of Early Learning Detection
(a) Random: mAP
(b) Random: mF1
(c) Dominant: mAP
(d) Dominant: mF1
Figure 5: Validation accuracies (mAP and mF1) over training epochs obtained with predefined warm-up lengths (10, 20, 30) and our proposed early learning detection strategy (AdaGC).
We evaluate the effectiveness of early learning detection for adaptively triggering GC by comparing AdaGC with variants that use manually predefined trigger points. Fig. 5 plots their validation accuracies (wrt GT labels) over training time. The proposed adaptive detection strategy successfully activates GC at an appropriate point, resulting in the most significant performance improvements. This demonstrates that the timing of GC activation is essential for model convergence stability. In contrast, triggering GC too early or too late leads to suboptimal results. Notably, activating GC too late leads to even worse performance, highlighting the detrimental effect of overfitting to label noise. Furthermore, we report their final test accuracies in Table VII. AdaGC achieves competitive or comparable mAP and mF1 scores compared to those obtained using fixed warm-up lengths, indicating the robustness of the proposed early learning detection strategy.
TABLE VII: Accuracies (%) on the Test Set Obtained With Different GC Trigger Settings After 70 Epochs
@@ IV-D1 Impact of Early Learning Detection
We demonstrate the necessity of using the teacher model for early learning detection by comparing the smoothness of noisy validation accuracies from the student and teacher models in Fig. 6. Although both models exhibit similar overall trends, the teacher model produces noticeably smoother accuracy curves. This stability makes it more suitable for reliable early learning detection. In contrast, detections based on the student model are more prone to local maxima, resulting in unstable and suboptimal outcomes.
@@
IV-D2 Effect of the Dual-EMA Mechanism
We investigate the effect of the proposed dual-EMA mechanism by comparing it with single-EMA and no-EMA settings, as shown in Table VIII. The dual-EMA strategy consistently achieves the most robust performance by striking a balance between stability and adaptability. In contrast, the single-EMA and no-EMA variants exhibit suboptimal results, particularly when the EMA smoothing on student predictions is removed. In both Random and Dominant noise settings, the EMA-smoothed student predictions are more stable than the teacher-only ones, owing to the teacher’s slower update rate. Nevertheless, without EMA smoothing (no-EMA), the student’s raw predictions tend to amplify errors, resulting in additional performance degradation. This trend is further confirmed in Fig. 7, where pseudo-label quality is shown to have a strong impact on AdaGC’s performance. As the pseudo-label quality deteriorates, performance drops markedly, especially for the teacher-only and no-EMA cases. The dual-EMA mechanism mitigates these issues by smoothing student predictions and integrating teacher outputs, thereby generating more reliable pseudo labels. Considering the varying update dynamics of the student and teacher models during training, a more adaptive combination strategy beyond the current half–half scheme is expected to enhance robustness, which we leave for future exploration.
TABLE VIII: Test set accuracies (%) obtained using different pseudo-label generation strategies, where T, S, and S denote teacher model predictions, EMA-smoothed student model predictions, and student model predictions without EMA, respectively
EMA Settings Random Dominant
@@ IV-D2 Effect of the Dual-EMA Mechanism
dual AdaGC 68.18 64.65 60.22 58.78
single Only T ($\gamma=1$) 61.74 63.49 42.01 46.40
Only S ($\gamma=0$) 67.93 61.52 60.11 57.42
T+S ($\gamma=0.5$) 53.78 61.94 27.50 35.71
no Only S ($\gamma=0$) 49.30 52.89 16.41 25.24
(a) Random: pseudo-label mAP
(b) Random: pseudo-label mF1
(c) Random: validation mAP
(d) Random: validation mF1
(e) Dominant: pseudo-label mAP
(f) Dominant: pseudo-label mF1
(g) Dominant: validation mAP
(h) Dominant: validation mF1
Figure 7: Pseudo-label quality (a), (b), (e), (f) and AdaGC validation accuracies (mAP and mF1) (c), (d), (g), (h) over training epochs using different pseudo-label generation strategies, where T, S, and S denote teacher model predictions, EMA-smoothed student model predictions, and student model predictions without EMA, respectively. (a)-(d) Random single-positive labels case. (e)-(h) Dominant single-positive labels case.
@@ IV-D3 Role of Mixup
We then investigate the effect of Mixup under both Random and Dominant noise settings. As shown in Table IX, removing Mixup consistently reduces performance, particularly in the Random case, with about a 2% drop in mAP and mF1. This indicates that Mixup enhances generalization by enforcing smoother decision boundaries. It can complement the GC mechanism in improving the robustness of AdaGC.
TABLE IX: Test accuracies (%) Obtained With and Without Mixup